For a certain study, the population proportion is known to be 0.8. If we draw a sample of 1000 units from that population what would be the 1 standard deviation?

A> 4%
B> 0.4%
C> 1.264%
D> None of the above
1
Expert's answer
2020-03-16T13:56:51-0400

Let "X" be a binomial random variable based on "n" trials with success probability "p."  

If the binomial probability histogram is not too skewed, "X" has approximately a normal distribution with "\mu=np" and "\sigma=\sqrt{p(1-p)/n}."

In practice, the approximation is adequate provided that both "np\geq10" and "n(1-p)\geq10," since there is then enough symmetry in the underlying binomial distribution.

Check normal approximation to binomial

"n=1000, p=0.8"

"np=1000(0.8)=800\geq10,"

"n(1-p)=1000(1-0.8)=200\geq10"

Hence we can use normal approximation to binomial. Then


"\sigma=\sqrt{p(1-p)/n}=\sqrt{0.8(1-0.8)/1000}\approx0.01265"

C> 1.264%